this is a fourth grade math problem. I need help. Total sales for the 9pm and 11pm shows were $602. Tickets for the 11pm show are $2 less than the 9pm show. How much were the ticket for eacj=h show?
9pm total tickets sold were 56 and 11pm total tickets sold were 42,
Please please help me
kammikam

Answer by stanbon(75887)


Total sales for the 9pm and 11pm shows were $602.
Tickets for the 11 pm show are $2 less than the 9 pm show. How much were the ticket for each show?
-----------------------------------------------------
Let the cost of tickets for the 9pm show be "x".
Then the cost of tickets for the 11pm show was "x-2"
---------------------------------------------------------
9pm total tickets sold were 56 and 11pm total tickets sold were 42,
Equations:
Value equation: 56x + 42(x-2) = 602
56x + 42x - 84 = 602
98x = 686
x = $7 (price of the 9 pm tickets)
x-2 = $5 (price of the 11 pm tickets)
--------------------------------------------
Cheers,
Stan H.